OpenStudy (anonymous):

Let f(x)=x^3-8(x). Calculate the difference quotient f(3+h)-f(3)/h for h=.1 h=.01 h=-0.1 h=-.1 If someone told you the derivative(slope of tangent line to the graph) of f(x) at x=3 was an integer, what would you expect it to be ? (I'm confused in how to approach this problem, I know I can't just plug in the given h's into the diff.Quot. so would I plug in f(x) into h ? )

OpenStudy (anonymous):

Would I start by pluging in my f(x) into the diff quotient ?

OpenStudy (anonymous):

Yes, but first you need to make the necessary substitutions. Notice the first: f(3+h) means to replace all x's in your original f(x) with 3+h. Same goes for f(3), where all x's would be replaced with 3 in your original f(x).

OpenStudy (anonymous):

Okay I'll try

OpenStudy (anonymous):

So [(3+h)^3 -8(3+h) - (x^3-8x)]/h ?

OpenStudy (anonymous):

The first part is correct. However, f(3) should just be a number, if f(3) is supposed to be there. f(3) = (3)^3 - 8(3) = 9-24 = -15

OpenStudy (anonymous):

That makes sense, but how I end up finding the values for the h's ?

OpenStudy (anonymous):

Well I don't think your question asks you to find h, it's asking you to substitute each value of h into the difference quotient one-by-one.

OpenStudy (anonymous):

Oh okay so I take the equation we came up with [(3+h)^3 -8(3+h) - (-15)]/h and plug in the h values right ?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

Oh okay let me see if I get it right :D thank you so much Pompeii00 !!!!!!:D

OpenStudy (anonymous):

Not a problem. Let me know how ya do.

OpenStudy (anonymous):

Just as clarification, notice f(3)=-15, make sure you incorporate that sign, so you'll actually have +15 in the difference quotient!!

OpenStudy (anonymous):

So far the values I have gotten @Pompeii00 turned out wrong :O for h=.1 I got 199.91 h=.01= 1819.0901

OpenStudy (anonymous):

What values are you expecting for this problem?

OpenStudy (anonymous):

I don't know its an online hw system :/

OpenStudy (anonymous):

Are you certain that the question, equation, and values of h are correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Nope, that's what the actual question looks like

OpenStudy (anonymous):

I've redone my math, I can't seem to figure out why the h values aren't working. At least from what I can see it makes sense. As for the derivative at x=3, i believe it to be 19. That's about it.

OpenStudy (anonymous):

Oh okay thanks for your help anyways !

OpenStudy (anonymous):

If anything comes to mind, I'll be sure to report it.

OpenStudy (anonymous):

Oh okay thanks, I'll keep searching :/